# Espresso: Robust Concept Filtering in Text-to-Image Models

### v2

@@ 6.2. Comparison with Filtering CRT
We compare with UD (Qu et al., 2023b), the state-of-the-art filtering CRT, across the three requirements and summarize the results. We use FNR to evaluate effectiveness on $\mathcal{D}_{te}^{u}$, FPR for utility on $\mathcal{D}_{te}^{a}$, and FNR for robustness on $\mathcal{D}_{adv}^{u}$.
@@
R1 Effectiveness. We report FNR across four concepts (nudity, violence, disturbing, and hateful). Espresso has better FNR for three of the four concepts: nudity, violence (in green), and hateful (blue for Espresso and red for UD). However, both Espresso and UD perform poorly on disturbing. We attribute this poor effectiveness on Group-1 concepts to the subjective description of $c^{u}$. Images for these concepts cover a wide variety of sub-concepts simultaneously which are not precisely identified for CRTs.
@@
Table 9. R1 (Effectiveness): Comparison with filtering CRT (UD) using FNR on unacceptable prompts (lower is better). We use red if FNR is $>$0.50; blue if FNR is between 0.25-0.50; green if FNR is $<$0.25.
@@
Table 10. R2 (Utility): Comparison with filtering CRT (UD) using FPR on acceptable prompts (lower is better). We use red if FPR is $>$0.50, blue if FPR is between 0.25-0.50; green if FPR is $<$0.25.
@@
R2 Utility. We present FPR in Table 10 across the four Group-1 concepts. As expected, we observe that both Espresso and UD perform well since they demonstrate a low FPR. This is expected since UD explicitly includes images containing $c^{a}$ while training the multi-headed classifier.
R3 Robustness. We report FNR on the dataset for adversarial prompts and corresponding images in Table 11. In addition to the four attacks from Table 8, recall that SneakyPrompt (Yang et al., 2024b) is specifically designed to attack filtering CRTs. Hence, we also include the evaluation against SneakyPrompt. Also, since CCE is not adaptive against filtering CRT’s, we evaluate UD and Espresso against CCE+. We are the first to show the effectiveness of different attacks against UD.
Table 11. R3 (Robustness): Comparison with filtering CRT (UD) using FNR on adversarial prompts (lower is better). We use red if FNR is $>$0.50; blue if FNR is between 0.25-0.50; and green if FNR is $<$0.25.
@@
We observe that Espresso is effective in thwarting $\texttt{p}^{adv}$s from PEZ+ while UD can be evaded. On all the remaining attacks, we show that Espresso is robust on more concepts than UD. As indicated before, all the Group-1 concepts are subjective and capture multiple sub-concepts. This ambiguity could be the reason for poor robustness on some of these concepts. Overall, Espresso shows better robustness than UD on the majority of the concepts across different attacks.
Figure 2. Espresso has a better trade-off across R1, R2, and R3 compared to filtering CRT (UD).
Summary. We present the trade-offs in the form of radar plots in Figure 2. We use (1- FNR) on $\mathcal{D}_{te}^{u}$ for R1, (1- FPR) on $\mathcal{D}_{te}^{a}$ for R2, and (1- FNR) on $\mathcal{D}_{adv}^{u}$ using CCE+ for R3. Hence, a higher value indicates better performance. Espresso has comparable utility to UD but outperforms in effectiveness and robustness. Overall, Espresso covers a larger area, and hence provides a better trade-off compared to UD.
7. Certifying Robustness of Espresso
We showed empirically that Espresso is robust against several state-of-the-art attacks (Section 6). Inspired by the literature on certified robustness against adversarial examples (Cohen et al., 2019), it is natural to ask whether a similar notion of certified robustness is possible for CRTs. None of the existing CRTs have considered certified robustness. To this end, we are the first to explore its feasibility for Espresso.
@@ 7. Certifying Robustness of Espresso
7.1. Theoretical Bound
Certified robustness aims to find provable guarantees that an ML model’s predictions (generally a classifier) are robust, i.e., the predictions do not change on adding noise to the input (Carlini et al., 2023b). Our goal is to have a similar provable robustness bound for a T2I model with Espresso. We want to find the maximum noise to an input which Espresso can tolerate.
We give advantage to $\mathcal{A}dv$ by assuming they can directly add adversarial noise to Espresso’s embeddings. This is a strong assumption as in practice, $\mathcal{A}dv$ can only send prompts to the T2I model. We revisit this assumption later in Section 7.3. Formally, given an unacceptable image $x^{u}$, $\mathcal{A}dv$ adds noise $\delta$ to its embeddings, $\phi_{x}(\texttt{$x^{u}$})$, such that $\texttt{F}(\phi_{x}(\texttt{$x^{u}$})+\delta)$ is misclassified as acceptable. Using this assumption, we specify the maximum noise $\delta$ added to the embeddings, $\phi_{x}(\texttt{$x^{u}$})$, that Espresso can tolerate in Theorem 7.1.
@@ 7.1. Theoretical Bound
Figure 3. Certified accuracy of Espresso vs. adversarial noise $\delta$, for a strong $\mathcal{A}dv$ with access to embeddings of generated images.
Theorem 7.1.
@@
(5) $F(\hat{x})=F(\hat{x}+\delta),\forall||\delta||\leq\left(1-\frac{\tau}{\tau+2|g(\hat{x})-\Gamma|}\right)||\hat{x}||.$
@@
Proof Sketch. We prove the above theorem by applying Lipschitz continuity over $g(\hat{x})$. F ($\cdot$) is the composition of the softmax function and the scaled cosine similarity over the embeddings, where both functions are Lipschitz continuous when $||\hat{x}||>0$. In the detailed proof, we compute the Lipschitz constant for the softmax function and scaled cosine similarity function respectively, which is 0.25 and $\frac{\tau}{||\hat{x}||}$. Then the Lipschitz constant for $g(\hat{x})$ will be $\frac{\tau}{2||x||}$ according to the chain rule. Finally, using the triangle inequality on $|g(\hat{x})-g(\hat{x}+\delta)|$, and plugging the bound in to the inequality, we get $g(\hat{x}+\delta)\geq\Gamma$.
@@
7.2. Empirical Validation
We now compute the maximum noise that Espresso can tolerate for each unacceptable image’s embedding using Equation 5. Following prior literature on certified robustness (Cohen et al., 2019), we compute the certified accuracy described in (Cohen et al., 2019) to evaluate the robustness of Espresso. Certified accuracy at radius $r$ is the fraction of unacceptable images which are correctly classified and are robust against adversarial noise $\delta>r$. This shows the robustness of Espresso against attacks under some noise $r$. A robust model will have a larger certified radius and higher certified accuracy. Since we add noise directly to $\phi_{x}(\texttt{$x^{u}$})$, we compare our certified accuracy with the accuracy of clean unacceptable images (without adversarial noise) which we refer as “clean accuracy”. Ideally, certified accuracy should be close to the accuracy of clean unacceptable images.
We present the results in Figure 3 for the three groups of concepts. Clean accuracy in Figure 3 is the certified accuracy at radius zero. Espresso is robust against $\delta<0.07$, incurring less than a $5\%$ drop in certified accuracy. When $\delta<0.15$, the certified accuracy remains higher than $50\%$ for all concepts. Espresso is particularly robust for some concepts in Group-2 (Grumpy Cat, R2D2, Captain Marvel), and Group-3 (Taylor Swift, Angelina Jolie, and Elon Musk). For these concepts, the certified accuracy remains the same for the clean unacceptable images until $\delta>0.15$. Further, Espresso is more robust for concepts where the clean accuracy is $1.00$ (CLIP accuracy from Table 6). We find that the robustness is higher for concepts on which Espresso is more accurate. We attribute this to the increased separation between acceptable and unacceptable concepts.
7.3. Practical Implications
Having discussed the theoretical bound and empirically validated it on different concepts, we now revisit the practicality of this bound. We first discuss the usefulness of the certification followed by our assumption about $\mathcal{A}dv$’s capability.
To illustrate this, we consider a concrete attack approach that $\mathcal{A}dv$ could adopt given its inability to directly add adversarial noise to embeddings: $\mathcal{A}dv$ can begin with unacceptable images and incorporate adversarial noise using standard evasion techniques (e.g., PGD (Madry et al., 2018)) to find an adversarial example that evades the Espresso classifier. $\mathcal{A}dv$ can then find the corresponding adversarial prompt using one of the attacks (PEZ+) in Section 2.3. We want to see if f still generates an adversarial image which evades Espresso. We use PGD to generate unacceptable images with adversarial noise, and PEZ+ to find their corresponding adversarial prompts. We find that f fails to generate an adversarial image which evades Espresso using the adversarial prompt. This is due to the adversarial-prompt-generation process being an approximation, which fails to fully capture all aspects of the adversarial image. Moreover, using the T2I model to generate the image from the adversarial prompt is unlikely to capture the adversarial noise due to the de-noising used in the diffusion model (Section 2.1). This claim is further supported by prior literature on the robustness of diffusion models (Kim et al., 2022; Chen et al., 2024; Zhang et al., 2023a; Xiao et al., 2022).
We compare the original adversarial images with the images generated from their adversarial prompts. We present one concept from each group in Appendix LABEL:app:adv: Figure LABEL:fig:adversarial. We find that the generated images are significantly different from the original adversarial images. This confirms our conjecture that the adversarial noise is not retained in the generated images. A more thorough exploration of such an attack is left as future work. Based on the above preliminary exploration, we conjecture that Espresso is likely to be robust against such attacks by $\mathcal{A}dv$ with realistic capabilities.
8. Related Work
Prompt Filters. We focus on image filters to detect $x^{u}$. Prior works have explored using prompt filters to identify unacceptable concepts before passing them to T2I models (Yang et al., 2024a). Such filters are being used for DALLE-2 (Ramesh et al., 2021), and MidJourney (Midjourney, 2024). However, these prompt filters are not robust and easy to evade (Ba et al., 2023; Ma et al., 2024). Hence, prior works have identified image filters are a better alternative, with the possibility of making it effective and robust (Li et al., 2024).
Additional CRTs. We use state-of-the-art CRTs for evaluation. Hence, we omit the discussion about prior CRTs which have been outperformed by those considered in this work (Schramowski et al., 2023; Prompt, [n.d.]; Radford et al., 2021). We discuss some recent fine-tuning CRTs.
Hong et al. (Hong et al., 2023) propose an approach to make minimal changes for concept removal. Instead of matching the entire distribution of features for $c^{u}$ to the distribution of $c^{a}$, they selectively modify the features with highest density in $c^{u}$. Wu et al. (Wu et al., 2024) propose concept removal as an unlearning problem which is solved as constraint optimization while preserving utility. Basu et al. (Basu et al., 2024) present a closed-form solution to remove concepts from $\phi$ by identifying causal features for one concept and overriding it with another concept. Since these approaches are similar to other fine-tuning CRTs considered in this work, and do not optimize for robustness, we conjecture that these approaches are susceptible to evasion.
Huang et al. (Huang et al., 2024) modify the cross entropy layers of $\epsilon_{\theta}$ to erase $c^{u}$ with adversarial training for better robustness against only two naïve attacks: Prompting4Debugging (Chin et al., 2023) and RingBell (Tsai et al., 2024). Li et al. (Li et al., 2024) modify the self-attention layers of $\epsilon_{\theta}$ using triplets of $x^{u}$, censored $x^{u}$, and $x^{a}$. They evaluate robustness against SneakyPrompt. However, the robustness for both these works against recent attacks which account for their CRTs is not clear.
Pham et al. (Pham et al., 2024b) fine-tune T2I model to generate a specific concept and subtract the task vector for this concept from the model parameters of the original model, thereby erasing $c^{u}$. They demonstrate robustness against CCE and RingBell but suffers from a trade-off between R2 and R3. Wu et al. (Wu et al., 2023) insert backdoors to textual inversion embeddings such that a $c^{u}$ results in a pre-defined image instead of $x^{u}$. Despite accounting for robustness, these CRTs incur trade-offs between R1, R2, and R3, inherent to fine-tuning CRTs.
Additional Attacks against CRTs. Prompting4Debugging uses PEZ while matching the noise between a T2I model with a CRT, and one without a CRT, during the reverse diffusion process. In our setting, T2I model with and without Espresso are identical, hence Prompting4Debugging reduces to PEZ. Rando et al. (Rando et al., 2022) propose PromptDilution against filters. However, as SneakyPrompt outperformed PromptDilution (Yang et al., 2024b), we do not consider it for evaluation.
[…]
9. Discussion and Future Work
We discuss efficiency of CRTs (Section 9.1), accounting for future attacks (Section 9.2), applicability/extensions of Espresso (Section 9.3), and a summary of this work (Section 9.4).
9.1. Efficiency of CRTs
@@
In Table 12, we report the execution time for fine-tuning or training the CRTs (average across ten runs). For related work, the configuration for fine-tuning/training is the same as specified by their respective paper to satisfy R1 and R2. These times were obtained from training on a single NVIDIA A100 GPU.
Table 12. Efficiency: CRT training time (mean across ten runs).
@@
CA (Kumari et al., 2023a) 60.03 $\pm$ 0.01 UCE (Gandikota et al., 2024) 0.24 $\pm$ 0.02
SA (Heng and Soh, 2023b) 95.10 $\pm$ 2.21 ESD (Gandikota et al., 2023a) 125.50 $\pm$ 0.00
SDD (Kim et al., 2023) 75.50 $\pm$ 3.21 UD (Qu et al., 2023b) 10.00 $\pm$ 2.03
FMN (Zhang et al., 2023b) 2.20 $\pm$ 0.01 Espresso 9.10 $\pm$ 0.05

### v4

@@ 6.2 Comparison with Filtering CRT
We now compare with UD [67], the state-of-the-art filtering CRT, across the three requirements and summarize the results. We use FNR to evaluate effectiveness on $\mathcal{D}_{te}^{u}$, FPR for utility on $\mathcal{D}_{te}^{a}$, and FNR for robustness on $\mathcal{D}_{adv}^{u}$.
@@
R1 Effectiveness. We report FNR across four concepts (nudity, violence, disturbing, and hateful). Espresso has better FNR for three of the four concepts: nudity, violence (in green), and hateful (blue for Espresso and red for UD). However, both Espresso and UD perform poorly on disturbing. We attribute this poor effectiveness on Group-1 concepts to the subjective description of $c^{u}$. Images for these concepts cover a wide variety of sub-concepts simultaneously which are not precisely identified for CRTs. Overall, Espresso is more effective than UD on most concepts.
@@
TABLE VIII: R1 (Effectiveness): Comparison with filtering CRT (UD) using FNR on unacceptable prompts (lower is better). We use red if FNR is $>$0.50; blue if FNR is between 0.25-0.50; green if FNR is $<$0.25.
@@
R2 Utility. We present FPR in Table IX across the four Group-1 concepts. As expected, we observe that both Espresso and UD have comparable utility they demonstrate a low FPR. This is expected since UD explicitly includes images containing $c^{a}$ while training the multi-headed classifier.
TABLE IX: R2 (Utility): Comparison with filtering CRT (UD) using FPR on acceptable prompts (lower is better). We use red if FPR is $>$0.50, blue if FPR is between 0.25-0.50; green if FPR is $<$0.25.
@@
R3 Robustness. We report FNR on the dataset for adversarial prompts and corresponding images in Table X. In addition to the four attacks from Table VII, recall that SneakyPrompt [66] is specifically designed to attack filtering CRTs. Hence, we also include the evaluation against SneakyPrompt. Also, since CCE is not adaptive against filtering CRT’s, we evaluate UD and Espresso against CCE+. We are the first to evaluate different attacks against UD.
TABLE X: R3 (Robustness): Comparison with filtering CRT (UD) using FNR on adversarial prompts (lower is better). We use red if FNR is $>$0.50; blue if FNR is between 0.25-0.50; and green if FNR is $<$0.25.
@@
We observe that Espresso is effective in thwarting $\texttt{p}^{adv}$s from PEZ+ while UD can be evaded. On all the remaining attacks, we show that Espresso is robust on more concepts than UD. As indicated before, all the Group-1 concepts are subjective and capture multiple sub-concepts. This ambiguity could be the reason for poor robustness on some of these concepts. Overall, Espresso is more robust than UD on majority of the concepts across different attacks.
Figure 3: Espresso has a better trade-off than filtering CRT (UD).
Summary. We present the trade-offs in the form of radar plots in Figure 3. We use (1- FNR) on $\mathcal{D}_{te}^{u}$ for R1, (1- FPR) on $\mathcal{D}_{te}^{a}$ for R2, and (1- FNR) on $\mathcal{D}_{adv}^{u}$ using CCE+ for R3. Hence, a higher value indicates better performance. Espresso has comparable utility to UD but outperforms in effectiveness and robustness. Overall, Espresso covers a larger area, and hence provides a better trade-off than UD.
7 Certifying Robustness of Espresso
We empirically showed that Espresso is robust against several state-of-the-art attacks (Section 6). Inspired by the literature on certified robustness against adversarial examples [12], it is natural to ask whether a similar notion of certified robustness is possible for CRTs. None of the existing CRTs have considered certified robustness. To this end, we are the first to explore its feasibility for Espresso.
@@ 7 Certifying Robustness of Espresso
7.1 Theoretical Bound
Certified robustness aims to find provable guarantees that an ML model’s predictions (generally a classifier) are robust, i.e., the predictions do not change on adding noise to the input [8]. Our goal is to have a similar provable robustness bound for a T2I model with Espresso. We want to find the maximum noise to an input which Espresso can tolerate.
We give advantage to $\mathcal{A}dv$ by assuming they can directly add adversarial noise to Espresso’s embeddings. This is a strong assumption as in practice, $\mathcal{A}dv$ can only send prompts to the T2I model. We revisit this assumption later in Section 7.3. Formally, given an unacceptable image $x^{u}$, $\mathcal{A}dv$ adds noise $\delta$ to its embeddings, $\phi_{x}(\texttt{$x^{u}$})$, such that $\texttt{F}(\phi_{x}(\texttt{$x^{u}$})+\delta)$ is misclassified as acceptable. Using this assumption, we specify the maximum noise $\delta$ added to the embeddings, $\phi_{x}(\texttt{$x^{u}$})$, that Espresso can tolerate in Theorem 1.
@@ 7.1 Theoretical Bound
Figure 4: Certified accuracy of Espresso vs. adversarial noise $\delta$, for a strong $\mathcal{A}dv$ with access to embeddings of generated images.
Theorem 1.
@@
$F(\hat{x})=F(\hat{x}+\delta),\forall||\delta||\leq\left(1-\frac{\tau}{\tau+2|g(\hat{x})-\Gamma|}\right)||\hat{x}||.$ (5)
@@
Proof Sketch. We prove the above theorem by applying Lipschitz continuity over $g(\hat{x})$. F ($\cdot$) is the composition of the softmax function and the scaled cosine similarity over the embeddings, where both functions are Lipschitz continuous when $||\hat{x}||>0$. In the detailed proof, we compute the Lipschitz constant for the softmax function and scaled cosine similarity function respectively, which is 0.25 and $\frac{\tau}{||\hat{x}||}$. Then the Lipschitz constant for $g(\hat{x})$ will be $\frac{\tau}{2||x||}$ according to the chain rule. Finally, using the triangle inequality on $|g(\hat{x})-g(\hat{x}+\delta)|$, and plugging the bound in to the inequality, we get $g(\hat{x}+\delta)\geq\Gamma$. We present the full proof in Appendix .6. We discuss extending the bound to prompt space in Section 9.1.
7.2 Empirical Validation
We now compute the maximum noise that Espresso can tolerate for each unacceptable image’s embedding using Equation 5. Following prior literature on certified robustness [12], we compute the certified accuracy described in [12] to evaluate the robustness of Espresso. Certified accuracy at radius $r$ is the fraction of unacceptable images which are correctly classified and are robust against adversarial noise $\delta>r$. This shows the robustness of Espresso against attacks under some noise $r$. A robust model will have a larger certified radius and higher certified accuracy. Since we add noise directly to $\phi_{x}(\texttt{$x^{u}$})$, we compare our certified accuracy with the accuracy of clean unacceptable images (without adversarial noise) which we refer as “clean accuracy”. Ideally, certified accuracy should be close to the accuracy of clean unacceptable images.
We present the results in Figure 4 for the three groups of concepts. Clean accuracy in Figure 4 is the certified accuracy at radius zero. Espresso is robust against $\delta<0.07$, incurring less than a $5\%$ drop in certified accuracy. When $\delta<0.15$, the certified accuracy remains higher than $50\%$ for all concepts. Espresso is particularly robust for some concepts in Group-2 (Grumpy Cat, R2D2, Captain Marvel), and Group-3 (Taylor Swift, Angelina Jolie, and Elon Musk). For these concepts, the certified accuracy remains the same for the clean unacceptable images until $\delta>0.15$. Further, Espresso is more robust for concepts where the clean accuracy is $1.00$ (CLIP accuracy from Table V). We find that the robustness is higher for concepts on which Espresso is more accurate. We attribute this to the increased separation between acceptable and unacceptable concepts.
7.3 Practical Implications
Having discussed the theoretical bound and empirically validated it on different concepts, we now revisit the practicality of this bound. We discuss the usefulness of the certification and revisit our assumption about $\mathcal{A}dv$’s capability.
To illustrate this, we consider a concrete attack that $\mathcal{A}dv$ could adopt given its inability to directly add adversarial noise to embeddings: $\mathcal{A}dv$ begins with unacceptable images and incorporate adversarial noise using standard evasion techniques (e.g., PGD [38]) to find an adversarial example that evades the Espresso classifier. $\mathcal{A}dv$ then finds the corresponding adversarial prompt using one of the attacks (PEZ+) in Section 2.3. We want to see if f still generates an adversarial image which evades Espresso. We use PGD to generate unacceptable images with adversarial noise, and PEZ+ to find their corresponding adversarial prompts. We find that f fails to generate an adversarial image which evades Espresso using the adversarial prompt. This is due to the adversarial-prompt-generation process being an approximation, which fails to fully capture all aspects of the adversarial image. Moreover, using the T2I model to generate the image from the adversarial prompt is unlikely to capture the adversarial noise due to the de-noising used in the diffusion model (Section 2.1). This claim is further supported by prior literature on the robustness of diffusion models [25, 9, 70, 63].
We compare the original adversarial images with the images generated from their adversarial prompts. We present one concept from each group in Appendix .7: Table XIII. We find that the generated images are significantly different from the original adversarial images. This confirms our conjecture that the adversarial noise is not retained in the generated images. A more thorough exploration of such an attack is left as future work. Based on the above preliminary exploration, we conjecture that Espresso is likely to be robust against such attacks by $\mathcal{A}dv$ with realistic capabilities.
8 Related Work
Prompt Filters. We focus on image filters to detect $x^{u}$. Prior works have explored using prompt filters to identify unacceptable concepts before passing them to T2I models [65]. Such filters are being used for DALLE-2 [47], and MidJourney [40]. However, these prompt filters are not robust and easy to evade [2, 37]. Hence, prior works have identified image filters are a better alternative, with the possibility of making it effective and robust [34].
Additional CRTs. We use state-of-the-art CRTs for evaluation. Hence, we omit the discussion about prior CRTs which have been outperformed by those considered in this work [52, 44, 46]. We discuss recent fine-tuning CRTs.
Hong et al. [23] propose an approach to make minimal changes for concept removal. Instead of matching the entire distribution of features for $c^{u}$ to the distribution of $c^{a}$, they selectively modify the features with highest density in $c^{u}$. Wu et al. [61] propose concept removal as an unlearning problem which is solved as constraint optimization while preserving utility. Basu et al. [3] present a closed-form solution to remove concepts from $\phi$ by identifying causal features for one concept and overriding it with another concept. Since these approaches are similar to other fine-tuning CRTs considered in this work, and do not optimize for robustness, making them susceptible to evasion.
Huang et al. [24] modify the cross entropy layers of $\epsilon_{\theta}$ to erase $c^{u}$ with adversarial training for better robustness against only two naïve attacks: Prompting4Debugging [11] and RingBell [56]. Li et al. [34] modify the self-attention layers of $\epsilon_{\theta}$ using triplets of $x^{u}$, censored $x^{u}$, and $x^{a}$. They evaluate robustness against SneakyPrompt. However, the robustness for both these works against recent attacks which account for their CRTs is not clear.
Pham et al. [43] fine-tune T2I model to generate a specific concept and subtract the task vector for this concept from the model parameters of the original model, thereby erasing $c^{u}$. They demonstrate robustness against CCE and RingBell but suffer from a trade-off between R2 and R3. Wu et al. [62] insert backdoors to textual inversion embeddings such that a $c^{u}$ results in a pre-defined image instead of $x^{u}$. Despite accounting for robustness, these CRTs incur trade-offs between R1, R2, and R3.
Additional Attacks against CRTs. Prompting4Debugging uses PEZ while matching the noise between a T2I model with a CRT, and one without a CRT, during the reverse diffusion process. In our setting, T2I model with and without Espresso are identical, hence Prompting4Debugging reduces to PEZ. Rando et al. [48] propose PromptDilution against filters. However, as SneakyPrompt outperformed PromptDilution [66], we do not consider it for evaluation.
[…]
9 Discussion and Future Work
We revisit our discussion on robustness (Section 9.1), applicability/extensions of Espresso (Section 9.2), and a summary of this work (Section 9.3).
9.1 Revisiting Robustness
Addressing Future Attacks. Recall that our certified robustness bound is loose (Section 7). We leave the improvement of the bound as future work.
Espresso maintains high robustness across all evaluated concepts. The design of new attacks which can preserve unacceptable concepts while evading Espresso is an open problem. Here, adversarial training for Espresso can be used to increase its robustness. We can fine-tune CLIP with this objective: $\mathcal{L}_{\text{Con-adv}}(\mathcal{D}_{adv}^{u})=$
[…]
Extending the Theoretical Bound to Prompts. To extend our theoretical bounds in Theorem 1 to the prompt space, we assume Lipschitz continuity, commonly used in ML [5], certified robustness [33], and more recently in the context of T2I models [10]. We consider the mapping from prompt space to embedding space is locally $L$-Lipschitz, i.e., the function $\phi_{x}(f(\text{{p}}))$ satisfies the following condition:
For each prompt p, there exists a $\delta_{\text{{p}}}>0$ such that $||\phi_{x}(f(\text{{p}}))-\phi_{x}(f(\text{{p}}+\delta_{\text{{p}}}))||\leq L||\delta_{\text{{p}}}||.$
Then, the inequality in Equation 5 can be related to the prompt space by a scalar factor $\frac{1}{L}$. Therefore, when $\|\delta_{\text{{p}}}\|\leq\frac{1}{L}\|\delta\|$, where $\|\delta\|$ satisfies inequality in Equation 5, the difference in embedding space will be bounded by $\|\delta\|$, and CLIP remains robust.
9.2 Applications and Extensions to Espresso
Filtering Multiple Concepts. Gandikota et. al. [19] have considered removing multiple concepts simultaneously. While we focus on filtering one concept at a time for comparison with other CRTs, Espresso can be extended by including multiple concepts simultaneously as well. Specifically, for $\texttt{F}(x,c^{u},c^{a})$ in Equation 1, instead of only specifying $c^{u}$ and $c^{a}$ for a single concept, we can include $c^{u}$ and $c^{a}$ for multiple concepts as a list. This is a simple configuration change with minimal fine-tuning or retraining in contrast to other filters (e.g., [67]) which require training with extensive datasets. Since none of the prior CRTs evaluate multiple $c^{u}$, we leave its evaluation as future work.
Filtering Artistic Styles. None of the prior filtering CRTs consider any copyrighted content and focus on only inapproriate Group-1 concepts. Espresso is the first filtering CRT which is applicable to concepts which are copyright-infringing (Group-2) or involve unauthorized use of personalized images (Group-3). Prior fine-tuning CRTs have considered removing artistic styles (e.g., painting in the style of Monet or Van Gogh) as copyrighted content [30, 16]. However, we observed that Espresso does not differentiate between images with and without artistic styles very well. Specifically, Monet painting as $c^{u}$ and painting as $c^{a}$, are very similar for Espresso, thus reducing its effectiveness. We leave the optimization of Espresso to account for artistic styles as future work.
9.3 Summary
Removing unacceptable concepts from T2I models is an important problem. However, none of the prior CRTs satisfy all three requirements simultaneously: effectiveness in preventing generation of images with unacceptable concepts, preserve utility on other concepts, and robustness against evasion. We propose Espresso, the first robust concept filtering CRT which provides a better trade-off among the three requirements compared to prior CRTs.
{arxiv}
Acknowledgments
This work is supported in part by Intel (in the context of Private AI consortium), and the Government of Ontario. Views expressed in the paper are those of the authors and do not necessarily reflect the position of the funding agencies.
F Function for Espresso classifier
$\alpha$ Regularization parameter for Espresso fine-tuning
$\overline{\cdot}$ Normalization function
$\epsilon_{\theta}$ Diffusion model parameterized by $\theta$
$\tau$ Temperature parameter
.5 Impact of Fine-tuning
We present the impact of fine-tuning of Espresso by reporting the effectiveness (CLIP accuracy) and utility (normalized CLIP score) for all the concepts in Table XII. We expect that fine-tuning improves utility on majority of the concepts (indicated with a higher normalized CLIP score) while maintaining the effectiveness (indicated with similar CLIP accuracy). We use red if effectiveness/utility degrades; blue if no significant change; green if effectiveness/utility is better.
TABLE XII: Benefit of Fine-Tuning (FT): On using fine-tuning, if effectiveness/utility degrades we use red; blue for no significant change; green if better. We measure CLIP accuracy for effectiveness (lower is better) and normalized CLIP score for utility (higher is better).
Concept Effectiveness Utility
w/o FT w/ FT w/o FT w/ FT
Nudity 0.15 $\pm$ 0.02 0.15 $\pm$ 0.06 0.83 $\pm$ 0.07 0.94 $\pm$ 0.08
Violence 0.21 $\pm$ 0.02 0.20 $\pm$ 0.05 0.55 $\pm$ 0.08 0.59 $\pm$ 0.11
Grumpy Cat 0.01 $\pm$ 0.01 0.00 $\pm$ 0.01 0.88 $\pm$ 0.03 0.98 $\pm$ 0.04
Nemo 0.12 $\pm$ 0.01 0.10 $\pm$ 0.02 0.24 $\pm$ 0.05 0.37 $\pm$ 0.02
Captain Marvel 0.04 $\pm$ 0.01 0.03 $\pm$ 0.01 0.36 $\pm$ 0.02 0.37 $\pm$ 0.03
Snoopy 0.09 $\pm$ 0.01 0.08 $\pm$ 0.02 0.64 $\pm$ 0.01 0.66 $\pm$ 0.03
R2D2 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.52 $\pm$ 0.03 0.66 $\pm$ 0.02
Taylor Swift 0.04 $\pm$ 0.01 0.02 $\pm$ 0.00 0.97 $\pm$ 0.01 0.98 $\pm$ 0.02
Angelina Jolie 0.12 $\pm$ 0.02 0.03 $\pm$ 0.00 0.97 $\pm$ 0.01 0.98 $\pm$ 0.03
Brad Pitt 0.02 $\pm$ 0.01 0.00 $\pm$ 0.00 0.98 $\pm$ 0.01 0.98 $\pm$ 0.01
Elon Musk 0.07 $\pm$ 0.01 0.03 $\pm$ 0.00 0.92 $\pm$ 0.02 0.97 $\pm$ 0.01
[…]
Nudity [Uncaptioned image] “artsy wbo venus moc bday oilandgoddess thru cropped endurindiefilm cropped r underetal <copyright sign>” [Uncaptioned image]
Nemo [Uncaptioned image] “moma fishy pet <heart emoji> constrafrm orange optimistic soaking …… vacancy trippy troubles groovy averages !̃” [Uncaptioned image]
Elon Musk [Uncaptioned image] “poet moderstare rested wakeupamerica (" blurred vaportide driverless <smiley emoji> broker celebrated mandelclap” [Uncaptioned image]
In Table XII, we observe that utility either remains the same or improves significantly for some concepts. Further, we observe that while the effectiveness remains the same for majority of the concepts, there is a benefit of fine-tuning to improve effectiveness for four concepts in Group-3 (i.e., Taylor Swift, Angelina Jolie, Brad Pitt, Elon Musk).
.6 Proof for Theorem 1
We now present the full proof for Theorem 1 below.
@@ F Function for Espresso classifier
.7 Retention of Adversarial Noise
We present the images generated from T2I models on passing adversarial prompts corresponding to adversarial examples in Figure XIII.
.8 Efficiency of CRTs
@@
In Table XIV, we report the execution time for fine-tuning or training the CRTs (average across ten runs). For related work, the configuration for fine-tuning/training is the same as specified by their respective paper to satisfy R1 and R2. These times were obtained from training on a single NVIDIA A100 GPU.
TABLE XIV: Efficiency: CRT training time (mean across ten runs).
@@
CA [30] 60.03 $\pm$ 0.01 UCE [19] 0.24 $\pm$ 0.02
SA [21] 95.10 $\pm$ 2.21 ESD [16] 125.50 $\pm$ 0.00
SDD [26] 75.50 $\pm$ 3.21 UD [67] 10.00 $\pm$ 2.03
FMN [68] 2.20 $\pm$ 0.01 Espresso 9.10 $\pm$ 0.05